OpenStudy (anonymous):

solve x - 3 divide by 4 = 1

OpenStudy (anonymous):

\[\frac{ x-3 }{ 4 } = 1\] x- 3 = 4 Can you solve it now?

OpenStudy (anonymous):

answer is 7 As, "4" from the denominator will go to the right hand side and multiply. x-3=1X4 x=4+3 (sign will change) x=7

OpenStudy (anonymous):

x-3